# Learning to Solve Multi-Objective Routing Problems on Multigraphs

### v1

@@ 5.1 Benchmarking Setup
Problems: In the asymmetric simple graph case we consider two standard routing problems: The Multi-Objective Traveling Salesman Problem (MOTSP) and Multi-Objective Capacitated Vehicle Routing Problem (MOCVRP). Specifically, we consider bi-objective variants of these. In the MOTSP, we sample two distances for each node pair. These are summarized over the tour to yield the objective values. Regarding the distances, we consider both the TMAT (Kwon et al. 2021) and XASY distributions (Gaile et al. 2022), where the former obeys the triangle inequality and the latter does not - making it more difficult. For the MOCVRP, we sample one distance according to these distributions. We then let the first objective be given by the total distance of a tour and the second objective by the makespan, which is the longest trip for a single vehicle. This is a standard benchmark problem in literature (Lin, Yang, and Zhang 2022).
In the multigraph setting, we consider the Multigraph MOTSP (MGMOTSP) with two objectives, where we sample two-dimensional edge distances using distributions called FIX and FLEX. Here, denotes the number of edges between each node pair. In the FLEX-distribution, we sample edge distances independently and remove edges which are dominated. In the FIX distribution we first sample edges independently, but then rearrange by sorting, so that the an edge that is best in one objective is the worst in the other objective and so on. Thus, in the FIX-distribution there are always exactly edges between each node pair, whereas FLEX has at most edges.
In Appendix C, we provide more in-depth explanations of the problems and distributions. We also provide results on Euclidean problems in Appendix E.
@@
Baselines: For the MOTSP and MGMOTSP, we utilize linear scalarization together with strong single-objective methods in the form of LKH3 (Helsgaun 2017) and Google OR-tools. We also benchmark against four Multi-Objective Evolutionary Algorithms (MOEAs): MOGLS (Ishibuchi and Murata 1998), MOEA/D (Zhang and Li 2007), NSGA-II (Deb et al. 2002) and NSGA-III (Blank, Deb, and Roy 2019). Below, we only show the results for the MOEA that performs best in each scenario. Finally, we design a neural baseline using the MatNet-architecture (Kwon et al. 2021) as encoder and a preference-conditioned MP network as decoder. We call this MatNet-Based Model (MBM) and our motivation is that this is a learning-based method that can be used almost ”off-the-shelf” by combining components from literature. In the multigraph setting, this architecture requires us to pre-process the input by pruning edges until we obtain a simple graph, which we do using linear scalarization.
In Appendix A and D, we provide further details about the neural benchmark and MOEAs respectively. In Appendix E, we also show results for all MOEAs as well as some other weaker heuristics.
@@
Model Settings: We train all models for epochs using randomly generated instances per epoch with the ADAM optimizer (Kingma and Ba 2015) and a learning-rate of . For GMS-EB we let the number of layers be , while GMS-DH has and . For both models we use as the embedding dimension and attention heads in the GREAT layers and MP decoders. Regarding the number of samples in the decoding, we set for the selection-decoder during training and during inference. In the routing-decoder, we always set to the problem size as customary in the POMO framework (Kwon et al. 2020). Finally, we augment the models during inference with a factor of . For MBM, this is done using the original MatNet-augmentation (Kwon et al. 2021), while for GMS we use the scaling-augmentation of Lischka et al. (2024).
@@
Metrics: We utilize the Hypervolume (HV) metric (Zitzler et al. 2003) to evaluate the performance of the methods. Formally, for a given set of points and a reference point that is dominated by all points in , the HV is defined as:
@@
where denotes the Lebesgue measure in , and represents the axis-aligned hyperrectangle bounded by and . Intuitively, a higher HV generally indicates better performance.
@@
In our experiments, we report the normalized HV, which is scaled to lie in the interval , averaged over 200 test instances. For the neural models we use 101 preferences during inference, which is also the number of subproblems for the benchmarks relying on scalarization. Apart from the HV, we provide the HV gap compared to the best-performing method as well as the inference time. In Appendix B, we also report some training times for the learning-based methods. For these methods, we used one NVIDIA Tesla A40 GPU with 48GB VRAM during training and inference. The non-learning-based methods were all tested using an Intel Xeon Gold 6338 CPU and 8 parallel processes.
@@ 5.2 Performance on Benchmark Problems
We show the performance of GMS together with the benchmarks in Table 1. As can be seen, GMS outperforms all benchmarks except LKH on the MOTSP and MGMOTSP, while it is the best performing model on the MOCVRP. This showcases that our architectures are generally applicable across a variety of distributions, instance sizes and problems. We note that the runtime for GMS-EB scales poorly with the problem size, but it is still fast compared to most benchmarks, especially without augmentation.
| FLEX2-20 | FLEX2-50 | |||||
| HV | Gap | Time | HV | Gap | Time | |
| MOEA/D | 0.75 | 9.19% | (2.7h) | 0.60 | 32.79% | (34h) |
| MBM | 0.72 | 12.89% | (4.9s) | 0.80 | 11.09% | (14s) |
| MBM (aug) | 0.73 | 11.98% | (34s) | 0.81 | 10.10% | (1.8m) |
| GMS-EB | 0.81 | 1.95% | (4.3s) | 0.89 | 0.97% | (60s) |
| GMS-DH | 0.80 | 3.37% | (3.1s) | 0.85 | 5.10% | (13s) |
| GMS-EB (aug) | 0.82 | 0.00% | (31s) | 0.90 | 0.00% | (7.9m) |
| GMS-DH (aug) | 0.82 | 0.33% | (16s) | 0.87 | 2.69% | (1.6m) |
5.3 Multigraph Problem without Clear Pruning
The main motivation behind learning to select edges in the multigraphs is to plan routes when no clear pruning method exists. Thus, we evaluate the performance on such a problem as well. Specifically, we consider an MGMOTSP with Time-Windows (MGMOTSPTW). Each edge is associated with a time and distance, while the objectives correspond to the number of violated time-windows and the total distance. Unlike the MGMOTSP with linear scalarization, it is not clear in this case which edges can be pruned apriori due to the more intricate time-window objective.
We display the performance on this problem in Table 2, and benchmark against MBM with LS pruning as well as MOEAs. Evidently, our models are the best performing ones, illustrating the value of learning to select edges in combination with learning-to-route.
@@ 5.4 Zero-Shot Generalization
We also test our models in a zero-shot setting on larger instances, which is displayed in Table 3. Both GMS-DH and GMS-EB remain competitive across both problems and all distributions.
@@
| GMS-DH | 0.95 | 2.27% | (2.2m) | 0.94 | 2.56% | (2.2m) | |
5.5 Ablation Study
Finally, we showcase the ability of the selection-decoder in GMS-DH to select beneficial edges. Table 4 shows the result when replacing this component with a simple pruning function that removes edges with suboptimal linearly scalarized cost. For fairness, we also benchmark using both linearly scalarized reward and Chebyshev reward. As discussed, in the linear case the pruning should be optimal.
Evidently, Chebyshev reward improves the results substantially. Moreover, learning the selection compared to using a simple function also seems slightly better, even in the simple case of the MGMOTSP.
| FLEX2-50 | ||
| HV | Gap | |
| GMS-DH (Ch. Reward) | 0.8862 | 1.84% |
| GMS-DH (Lin. Reward) | 0.8781 | 2.74% |
| GMS-DH Simple (Lin. Reward) | 0.8789 | 2.65% |
| GMS-DH Simple (Ch. Reward) | 0.8847 | 2.00% |
6 Conclusion and Future Work
In this paper, we propose two methods to learn to construct solutions for routing problems on multigraphs. We highlight that while we solve problems with multiple objectives, our models can be valuable in all scenarios with no clear choice between alternative paths. In future works, we will apply our methods on more advanced problems.
A key weakness of our approach is that GMS-DH does not quite reach the same performance as GMS-EB in most cases. In this regard, there are works that are tangential to ours that could be utilized to obtain even better results for either GMS-DH or GMS-EB. Examples include the HV training of Lu et al. (2024) and diversity enhancement of Chen et al. (2023b). In the future, we will also explore decomposition methods like Ye et al. (2024) to extend our approach to even larger problem sizes.
Finally, from a broader perspective, we want to stress that routing problems on multigraphs are not well explored in literature. We believe one major reason is because of the complexity this formulation introduces. Learning-based methods can help alleviate this weakness and make the multigraph formulation more appealing.
@@ Acknowledgements
Appendix A Architecture Details
@@
A.1 Edge-Based Decoder in GMS-EB
In GMS-EB, we utilize an edge-based multi-pointer decoder which is based on Jin et al. (2023). Its mathematical formulation is the following. Given trainable weight matrices we form the query, representing the current routing context, as
@@ A.1 Edge-Based Decoder in GMS-EB
It must be noted that in our case, we preference-condition the decoder using a hyper-network. This is done by letting a network generate the weights and .
A.2 Selection-Decoder in GMS-DH
The task of this decoder is to prune parallel edges in the intermediate stage of GMS-DH, leaving only one between each node pair. Initially, we apply GREAT NB layers on the edge embeddings from the encoder to obtain embeddings specialized for this task. In the experiments we set . These layers are not preference conditioned. Hence, they do not need to be rerun for each preference. Next, given two nodes, we want to compute the probability of retaining each (directed) edge in between. For this purpose, we essentially utilize a multi-pointer score calculation similar to (6). Given a node pair , and the set of edges in between, , let the query be the average embedding according to
@@ A.2 Selection-Decoder in GMS-DH
A.3 MatNet-Based Benchmark Model
We visualize this architecture in Figure 3. The encoder consists of MatNet-layers (Kwon et al. 2021), while the decoder consists of an MP decoder (Jin et al. 2023) preference-conditioned with a hyper-network. The MatNet-architecture operates on bipartite graphs. It utilizes mixed-score attention layers together with fully-connected layers and normalization as in standard transformer layers (Vaswani et al. 2017). These attention layers ”mix” internally computed attention scores with external information from e.g. a distance matrix . To accommodate the use of several distance matrices, in our multi-objective setting, we follow the idea in Appendix A of the original paper. That is, we mix the internal attention score with all distance matrices using an MLP with inputs.
@@ A.3 MatNet-Based Benchmark Model
Lastly, we note that MatNet is not well-defined on multigraphs due to the fact that it utilizes the distance matrix to represent unique distances between nodes. It is possible to cast multigraphs into simple graphs, but not without increasing the number of nodes and edges drastically. Thus, we must pre-prune a multigraph before passing it through the encoder. This can be done with some sort of heuristic in the general case, but, as discussed, under LS and considering the MOTSP, a simple pruning method is to remove edges with sub-optimal linearly scalarized cost. Moreover, it can be noted that the encoder needs to be rerun for each new preference, which results in a slower architecture than if the encoding is done once.
Appendix B Model Training

### v2

@@ 5.1 Benchmarking Setup
Problems
In the asymmetric simple graph case we consider two standard routing problems: The Multi-Objective Traveling Salesman Problem (MOTSP) and Multi-Objective Capacitated Vehicle Routing Problem (MOCVRP). Specifically, we consider bi-objective variants of these. In the MOTSP, we sample two distances for each node pair. These are summarized over the tour to yield the objective values. Regarding the distances, we consider both the TMAT (Kwon et al. 2021) and XASY distributions (Gaile et al. 2022), where the former obeys the triangle inequality and the latter does not - making it more difficult. For the MOCVRP, we sample one distance according to these distributions. We then let the first objective be given by the total distance of a tour and the second objective by the makespan, which is the longest trip for a single vehicle. This is a standard benchmark problem in literature (Lin, Yang, and Zhang 2022).
In the multigraph setting, we consider the Multigraph MOTSP (MGMOTSP) with two objectives, where we sample two-dimensional edge distances using distributions called FIX and FLEX. Here, denotes the number of edges between each node pair. In the FLEX-distribution, we sample edge distances independently and remove edges which are dominated. In the FIX distribution we first sample edges independently, but then rearrange by sorting, so that the an edge that is best in one objective is the worst in the other objective. Thus, in the FIX-distribution there are always exactly edges between each node pair, whereas FLEX has at most edges.
In Appendix D, we provide more in-depth explanations of the problems and distributions. We also provide results on Euclidean problems in Appendix G.
@@
For the MOTSP and MGMOTSP, we utilize linear scalarization together with strong single-objective methods in the form of LKH3 (Helsgaun 2017) and Google OR-tools (Perron and Furnon 2019). We also benchmark against four Multi-Objective Evolutionary Algorithms (MOEAs): MOGLS (Ishibuchi and Murata 1998), MOEA/D (Zhang and Li 2007), NSGA-II (Deb et al. 2002) and NSGA-III (Blank, Deb, and Roy 2019). Below, we only show the results for the MOEA that performs best in each scenario.
Finally, we also design a neural baseline based on the MatNet-architecture (Kwon et al. 2021) as encoder and a preference-conditioned MP network as decoder. We call this MatNet-Based Model (MBM) and our motivation is that it is a method that can be used almost ”off-the-shelf” by combining components from literature. In the multigraph setting, this architecture requires us to pre-process the input by pruning edges until we obtain a simple graph, which we do using linear scalarization following Proposition 1.
Model Settings
@@
We train all models for epochs using randomly generated instances per epoch with the ADAM optimizer (Kingma and Ba 2015). We use the batch size and the learning-rate . For GMS-EB we let the number of layers be , while GMS-DH has and . For both models we set the embedding dimension to 128 and use attention heads in the GREAT layers and MP decoders. Regarding the number of samples in the decoding, we set for the selection-decoder during training and during inference. In the routing-decoder, we always set to the problem size as customary in the POMO framework (Kwon et al. 2020). Finally, we augment the models during inference with a factor of . For MBM, this is done using the original MatNet-augmentation (Kwon et al. 2021), while for GMS we use the scaling-augmentation of Lischka et al. (2025).
Metrics
@@
We utilize the Hypervolume (HV) metric (Zitzler et al. 2003) to evaluate the performance of the methods. Formally, for a given set of points and a reference point that is dominated by all points in , the HV is defined as:
@@
where denotes the Lebesgue measure in , and represents the axis-aligned hyperrectangle bounded by and . Intuitively, a higher HV generally indicates better performance, since the points in the obtained pareto front are further from the reference.
@@
We report the normalized HV, which is scaled to lie in the interval , averaged over 200 test instances. For the neural models we use 101 preferences during inference, which is also the number of subproblems for the benchmarks relying on scalarization. Apart from the HV, we provide the HV gap compared to the best-performing method as well as the inference time. In Appendix C, we also report some training times for the learning-based methods.
Hardware
We conducted training and inference for the learning-based methods using a single NVIDIA Tesla A40 GPU with 48GB of VRAM. For the non-learning-based methods, evaluations were performed on an Intel Xeon Gold 6338 CPU and, to enable a more fair comparison, we report execution time based on solving multiple instances in parallel. Akin to Kwon et al. (2021), we set the number of parallel processes to 8.
@@ 5.2 Performance on Benchmark Problems
We show the performance of GMS together with the other methods in Table 1. Both variants of GMS outperform most benchmarks on all problems. On the MOCVRP they are the best performing methods, while only LKH outperform them on the MOTSP and MGMOTSP. This showcases that our architectures are generally applicable across a variety of distributions, instance sizes and problems. As for the baseline neural method, MBM, it performs quite well across many cases, especially on the MGMOTSP, but severely under performs on some distributions, e.g., XASY100 for the MOCVRP.
However, MBM is usually slightly faster than GMS-DH, owing to the MatNet architecture being faster and more memory-efficient than GREAT (for a comparison, see the results by Lischka et al. (2025)). GMS-DH has a small advantage for the MGMOTSP with nodes due to only encoding each multigraph once, but this is offset again for nodes. Nonetheless, both GMS models remain efficient relative to the other benchmarks, despite the slower scaling of GMS-EB. In particular, GMS-DH as well as GMS-EB without augmentation are significantly faster than LKH.
Apart from the previously mentioned benchmarks, we also evaluate a variant of GMS-DH with Pre-Pruning (GMS-DH PP) applied before the encoder in the MGMOTSP setting. This pre-pruning is the same as for MBM, and the selection head is removed accordingly.
The poor performance of GMS-DH PP in Table 1 highlights the importance of explicitly designing for the multigraph structure, even when a theoretically sound pruning strategy is available. Since GMS-DH PP requires re-running the encoder for each preference, it is approximately 3 slower than GMS-DH. We also hypothesize that the pre-pruning step shifts the data distribution in a way that GMS-DH is ill-equipped to model, resulting in performance that is worse than MBM with the same pruning strategy. Indeed, in our experience GMS-DH PP tends to produce Pareto fronts with poor diversity, clustered around central preference regions
@@ 5.3 Zero-Shot Generalization
We also test our models on larger instances not seen during training, and report the results in Table 2. Both GMS-DH and GMS-EB remain competitive across both problems and all distributions in this zero-shot setting.
@@
GMS-DH | 0.95 | 2.27% | (2.2m) | 0.94 | 2.56% | (2.2m) |
5.4 More Difficult Multigraph Problem
Finally, we revisit one of the main motivations behind this work: routing for problems without clear a priori edge selection. Thus, we evaluate the performance on a more complex problem. Specifically, inspired by Ben Ticha et al. (2017), we consider an MGMOTSP with Time-Windows (MGMOTSPTW). Each edge is associated with a time and distance, while the objectives correspond to the number of violated time-windows and the total distance. Remark that, unlike the MGMOTSP, it is not clear in this case which edges can be pruned due to the more intricate time-window objective.
Besides MBM with linear pre-pruning and MOEAs, we benchmark against a simplified version of GMS-DH in order to ablate its ability to learn which edges are optimal. In this simplified version, we replace the selection-decoder with a simple function that removes edges with suboptimal linear cost. Compared to pre-pruning the multigraph before the encoder, this approach is smarter, as it removes the need to reencode. Furthermore, this simple variant actually performs quite well on the MGMOTSP, which we show and discuss in Appendix F.
We display the results for MGMOTSPTW in Table 3. Our two models are the best performing ones by quite a large margin. Notably, vanilla GMS-DH also outperforms the simplified variant, which illustrates the value of learning to select beneficial edges in comparison to using a simple pruning heuristic.
FLEX2-50 | FIX2-50 | |||||
HV | Gap | Time | HV | Gap | Time | |
MOEA/D | 0.60 | 32.79% | (34h) | 0.60 | 35.73% | (36h) |
MBM | 0.80 | 11.09% | (14s) | 0.64 | 31.18% | (14s) |
MBM (aug) | 0.81 | 10.10% | (1.8m) | 0.66 | 28.58% | (1.9m) |
GMS-EB | 0.89 | 0.97% | (60s) | 0.93 | 0.55% | (59s) |
GMS-DH | 0.85 | 5.10% | (13s) | 0.91 | 1.74% | (12s) |
GMS-EB (aug) | 0.90 | 0.00% | (7.9m) | 0.93 | 0.00% | (7.9m) |
GMS-DH (aug) | 0.87 | 2.69% | (1.6m) | 0.92 | 0.78% | (1.7m) |
GMS-DH Simple | 0.83 | 7.01% | (13s) | 0.85 | 8.93% | (12s) |
6 Conclusion
In this paper, we propose two learning-based approaches for solving multi-objective routing problems, such as the MOTSP, on multigraphs and asymmetric graphs. Both methods utilize graph neural networks for encoding the graph structure but differ in their route construction strategies. The first method is edge-based and directly builds tours on the multigraph, while the second employs a two-stage decoding process, consisting of non-autoregressive pruning of the multigraph followed by route construction.
Empirical results show that both methods perform well across a variety of settings. While the edge-based model is simpler and usually demonstrates slightly superior performance, the second model is more scalable due to lower memory and runtime demands.
Finally, we highlight that our models can be valuable in all routing scenarios where multigraphs are appropriate, i.e., when there are multiple competing edges between each node pair. In future work, we plan to use our setup to address single-objective problems with hard constraints as well as stochastic variants.
@@ Acknowledgements
Appendix A Extended Related Work
In this section, we complement the literature review in the main text with a more thorough discussion of related work.
A.1 Single-Objective Learning for Routing
In the autoregressive paradigm, Vinyals, Fortunato, and Jaitly (2015) were early to propose learning solutions to the TSP. They utilized an oracle solver for supervised learning, and Bello et al. (2017) later introduced reinforcement learning for training. By using attention models and inherent symmetries in the problem, first Kool, van Hoof, and Welling (2018) and then Kwon et al. (2020) obtained greatly improved optimality gaps on various routing problems. Regarding asymmetric problems, Kwon et al. (2021) proposed MatNet, a mixed-attention architecture working on bi-partite graphs. More recent works are Jin et al. (2023); Drakulic et al. (2023); Chen et al. (2024); Bi et al. (2024). These papers have improved results, can handle constraints better and can scale to even larger problems than previously. Another recent notable contribution was made by Drakulic, Michel, and Andreoli (2025), who proposed a jointly trained generalist model to solve a multitude of combinatorial optimization problems, including examples from routing.
In the non-autoregressive category, some works are Joshi, Laurent, and Bresson (2019); Fu, Qiu, and Zha (2021); Qiu, Sun, and Yang (2022); Sun and Yang (2023). Typically, these methods require more sophisticated procedures than simple sampling in the decoding, such as Monte–Carlo tree search, to perform competitively, but scale better to large instances due to lower space and time complexity. Thus, an interesting recent work combining autoregressive and non-autoregressive construction to solve large instances (up to nodes for the TSP) while maintaining performance is Ye et al. (2024).
A.2 Multi-Objective Learning for Routing
Early works treating MO routing using learning utilize multiple models, where one model is responsible for one preference between the objectives (Li, Zhang, and Wang 2021; Wu, Wang, and Zhang 2020). This quickly becomes infeasible as the number of models required grows exponentially with the number of objectives (Ehrgott 2005). Pareto-set learning for multi-objective routing using a single model was proposed by Lin, Yang, and Zhang (2022). They used the hyper-networks of Navon et al. (2021) to condition the decoder on the preference. Recent studies that feature improved results with a similar conditioning method include Fan et al. (2025) and Lu et al. (2024). An alternative approach is to use meta-learning to train a single model that can be quickly fine-tuned over a few steps to solve individual subproblems. Two representative works in this direction are Zhang et al. (2021) and Chen et al. (2023a).
Another significant recent work is that of Chen et al. (2023b), in which two diversity enhancement methods are introduced. The authors combine these methods with both the meta-learning approach and hyper-network approach to achieve improved results.
Besides tackling asymmetric problems, our work is novel in the MO routing setting as most methods utilize autoregressive construction, while one of our models combine autoregressive and non-autoregressive construction.
A.3 Routing on Multigraphs
Successful approaches to tackle MO shortest path problems on multigraphs include genetic algorithms (Beke, Weiszer, and Chen 2021), A∗-related methods (Weiszer, Burke, and Chen 2020) and hybrid approaches combining A∗ with learning (Liu, Chen, and Weiszer 2022). These methods all aim to find a Pareto set of shortest paths between two destinations.
For combinatorial routing problems such as VRPs on multigraphs, heuristics and exact solvers have been proposed. The first study in this regard was conducted by Garaix et al. (2010), who designed a simple heuristic and branch-and-price method for a single-objective on-demand transportation problem. Other works for single-objective constrained problems or multi-objective problems are Lai, Caliskan Demirag, and Leung (2016); Ben Ticha et al. (2019); Tikani, Setak, and Demir (2021). In addition, Ben Ticha et al. (2017) clearly concluded that a multigraph setup yields significantly lower costs compared to a simple graph representation for the capacitated VRP with time-windows. At the same time, the computation time increases significantly under this perspective, and the authors highlight the need for new heuristics.
Appendix B Architecture Details
@@
B.1 Edge-Based Decoder in GMS-EB
In GMS-EB, we utilize an edge-based multi-pointer decoder which is inspired by Jin et al. (2023). Its mathematical formulation is the following. Given trainable weight matrices we form the query, representing the current routing context, as
@@ B.1 Edge-Based Decoder in GMS-EB
In our case, we also preference-condition the decoder using a hyper-network. This is done by letting a multi-layer perceptron generate the weights and .
B.2 Selection-Decoder in GMS-DH
The task of this decoder is to prune parallel edges in the intermediate stage of GMS-DH, leaving only one between each node pair. Initially, we apply GREAT NB layers on the edge embeddings from the encoder to obtain embeddings specialized for this task. In the experiments we set . These layers are not preference conditioned and hence they do not need to be rerun for each preference. Next, given two nodes, we want to compute the probability of retaining each (directed) edge in between. For this purpose, we essentially utilize a multi-pointer score calculation similar to (23). Given a node pair , and the set of edges in between, , let the query be the average embedding according to
@@ B.2 Selection-Decoder in GMS-DH
B.3 MatNet-Based Benchmark Model
We visualize this architecture in Figure 3. The encoder consists of MatNet-layers (Kwon et al. 2021), while the decoder consists of an MP decoder (Jin et al. 2023) preference-conditioned with a hyper-network.
The MatNet-architecture operates on bipartite graphs. It utilizes mixed-score attention layers together with fully-connected layers and normalization as in standard transformer layers (Vaswani et al. 2017). These attention layers ”mix” internally computed attention scores with external information from e.g. a distance matrix . To accommodate the use of several distance matrices, , in our multi-objective setting, we follow the idea in Appendix A of the original paper. That is, we mix the internal attention score with all distance matrices using an MLP with inputs.
@@ B.3 MatNet-Based Benchmark Model
Lastly, we note that MatNet is not well-defined on multigraphs due to the fact that it utilizes the distance matrix to represent unique distances between nodes. It is possible to cast multigraphs into simple graphs, but not without increasing the number of nodes and edges drastically. Thus, we must pre-prune a multigraph before passing it through the encoder. This can be done with some sort of heuristic in the general case, but under linear scalarization and considering the MOTSP, an optimal pruning method is to remove edges with suboptimal scalarized cost. We utilize this pruning method for both multigraph problems. Moreover, it can be noted that the encoder needs to be rerun for each new preference, which results in a slower architecture than if the encoding is done once.
Appendix C Model Training